Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

From a box with 5 green, 5 red, 7 black and 3 yellow marbles are drawn 7 marbles at random without replacement. Find the following probabilities: P(2red,2green,2black,1yellow) P(the number of black marbles is 5) When the seven marbles are drawn at random with replacement find P(2red,2green,2black,1yellow) P(the number of black marbles is 5) P(The number of black marbles is amaller than 4)

OpenStudy (anonymous):

First find the prob that ALL of them ARE red, which is (6/11)(5/10)(4/9) = 2/33 So the prob that not all of them are red = 1-2/33 =31/33

OpenStudy (anonymous):

first one is \[\frac{\dbinom{5}{2}\times \dbinom{5}{2}\times \dbinom{7}{2}\times \dbinom{3}{1}}{\dbinom{20}{7}}\] do you know how to compute these?

OpenStudy (anonymous):

yeahh its like |dw:1334661870436:dw|

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

second one is \[\frac{\dbinom{7}{5}\times \dbinom{13}{2}}{\dbinom{20}{7}}\] yes same think just different notation

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

\[\dbinom{20}{7}=77520\] yw

OpenStudy (anonymous):

yeas yeas :)

OpenStudy (anonymous):

do you know how to do the ones with replacement? :)

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!